# Computing Wasserstein Barycenters through Gradient Flows

### v2

@@ Abstract
Wasserstein barycenters provide a powerful tool for aggregating probability measures, while leveraging the geometry of their ambient space. Existing discrete methods suffer from poor scalability, as they require access to the complete set of samples from input measures. We address this issue by recasting the original barycenter problem as a gradient flow in the Wasserstein space. Our approach offers two advantages. First, we achieve scalability by sampling mini-batches from the input measures. Second, we incorporate functionals over probability measures, which regularize the barycenter problem through internal, potential, and interaction energies. We present two algorithms for empirical and Gaussian mixture measures, providing convergence guarantees under the Polyak-Łojasiewicz inequality. Experimental validation on toy datasets and domain adaptation benchmarks show that our methods outperform previous discrete and neural net-based methods for computing Wasserstein barycenters.
@@ 1 Introduction
(a) Static barycenter.
(b) Barycenter flow.
Figure 1: In (a), we show the usual static notion of the Wasserstein barycenter (in red), which minimizes the sum of distances to the input measures (in blue). In (b), we show our notion of barycenter as a gradient flow, flowing an initial measure $P_{0}$ (purple) to the barycenter $P^{\star}$ (yellow) of the input measures.
Wasserstein barycenters [agueh2011barycenters] constitute a building block in the analysis of probability measures under Optimal Transport (OT) [villani2008optimal]. This mathematical object extends the notion of the barycenter of points to Wasserstein spaces, thus defining a notion of the average of probability measures. Since it relies on the Wasserstein metric, these barycenters capture the geometry of the underlying space over which the probability measures are defined [kloeckner2010geometric]. As such, they contributed to several areas of machine learning, including model averaging [singh2020model], ensembling [coz2023barycenter], data augmentation [zhu2023interpolation], distillation [liu2023dataset, montesuma2024multisource], dictionary learning [schmitz2018wasserstein], domain adaptation [montesuma2021wasserstein, montesuma2023multi, montesuma2024lighter], and Bayesian learning [srivastava2018scalable].
Existing methods for computing Wasserstein barycenters can be divided into three categories. The first category relies on discretization of the input measures and the barycentric measure through empirical measures. These algorithms may be subdivided into fixed-support [cuturi2014fast, Algorithm 1][benamou2015iterative, janati2020debiased], and free-support [cuturi2014fast, Algorithm 2][montesuma2023multi].
The second category relies on parametric models for the input measures, such as Gaussian, or Gaussian mixture models. In the Gaussian case, these barycenters are known as Bures-Wasserstein barycenters [chewi2020gradient, kroshnin2021statistical]. For Gaussian mixtures, [delon2020wasserstein] shows connections with multi-marginal OT [pass2015multi] between the components of the input mixtures, while [montesuma2024lighter] shows that, for components with diagonal covariances, there is an iterative algorithm similar to [cuturi2014fast, Algorithm 2].
The third category of methods for computing Wasserstein barycenters relies on neural nets and comes from the effort to scale OT to measures in high-dimensional spaces with a large number of samples. In this sense, [korotin2021continuous] proposes parameterizing the barycentric measure through a specific kind of neural net architecture, known as Input Convex Neural Networks (ICNNs) [amos2017input]. Furthermore, [kolesov2024estimating] proposes a bi-level adversarial learning approach that can be used for general ground costs, which is made more robust by semi-unbalanced OT [gazdieva2024robust]. More recently, [visentin2025computing] proposes computing barycenters with normalizing flows.
Current Wasserstein barycenter algorithms mostly focus on scalability with respect to the number of samples $n$, and dimensions $d$. In this sense, neural networks have a clear advantage over discrete methods, since they can operate at the level of mini-batches. There is another level of scalability, oftentimes overlooked, which is with respect the number of input measures $K$. So far, only [visentin2025computing] has considered experiments in this context. In this regard, methods such as [korotin2021continuous, korotin2022wasserstein] are difficult to scale, as they require $\mathcal{O}(K)$ neural networks to compute the Wasserstein barycenter.
In this paper we propose a new perspective on discrete Wasserstein barycenters, through the lens of Wasserstein Gradient Flows (WGFs) [ambrosio2008gradient, santambrogio2017euclidean], which is an appealing framework for optimization in the space of probability measures. This new understanding of Wasserstein barycenters has several advantages. First, we can scale the discrete barycenter problem, by drawing mini-batches from the input measures. Second, we can scale with respect to the number of input measures, as the representation of the barycentric measure is independent of $K$. Third, based on the established literature on gradient flows, we can define regularized Wasserstein barycenters with suitable functionals on the space of probability measures, which were not possible with previous methods. Lastly, we can modify the underlying metric to integrate label information between samples, leading to barycenters that respect class structure and provide more accurate and semantically meaningful estimates. We show an overview of our approach in Figure 1. Our contributions are:
  • We introduce a unified framework for computing Wasserstein barycenters based on gradient flows, which provides a principled and flexible approach to optimizing in the space of probability measures.
  • We generalize algorithms from the state-of-the-art [cuturi2014fast, montesuma2023multi, montesuma2024lighter] under a common framework.
  • We offer theoretical guarantees for the convergence of our algorithms, which also apply to previous works.
  • We offer a new theoretical result for the mixture-Wasserstein distance between labeled Gaussian mixtures [delon2020wasserstein, montesuma2024lighter].
  • We conduct extensive experiments benchmarking existing methods, including neural network-based Wasserstein barycenter algorithms [korotin2021continuous, kolesov2024estimating, gazdieva2024robust, visentin2025computing], on both toy datasets and domain adaptation  [^1] In contrast with discrete methods, most experiments in neural network solvers involve generative modeling. We thus establish a bridge between these methods and domain adaptation benchmarks, and show that our approach establishes a new state-of-the-art.
Paper organization. Section 2 presents the notation and preliminaries of our paper. Section 3 covers our proposed algorithms, including the theoretical results. Section 4 describes our experiments. Finally, section 5 concludes this paper.
@@ 2 Background
Throughout this paper, $(\Omega,d)$ is a metric space. We denote by $P$ and $Q$, probability measures on $\mathcal{P}(\Omega)$, which is the space of measures over $\Omega$. For $p\in[1,+\infty)$ we further denote by $\mathcal{P}_{p,ac}(\Omega)$, the sub-set of absolutely continuous measures w.r.t. the Lebesgue measure such that $\int_{\Omega}d(z,z_{0})^{p}dP(z)<+\infty$. In our case, we have access to these measures through their samples, denoted $z^{(P)}\sim P$. The set $\Delta_{K}$ denotes the $K$-simplex, i.e., $y\in\Delta_{n}=\{y\in\mathbb{R}^{n}:\sum_{i=1}^{n}y_{i}=1\text{ and }y_{i}\geq 0\text{ }\forall i\}$. We denote functionals over $\mathcal{P}_{2,\text{ac}}$ with blackboard bold letters (e.g., $\mathbb{F}$), and their (Wasserstein) gradients by $\nabla\mkern-10.0mu\nabla$. We offer a short introduction in Section 3 in the Appendix.
2.1 Empirical Optimal Transport
In this section, given $n$ i.i.d. samples from $P$, we can approximate a probability $P$ empirically through,
$\hat{P}(z)=\dfrac{1}{n}\sum_{i=1}^{n}\delta(z-z_{i}^{(P)}),$ (1)
which gives a finite parametrization for $P$. Henceforth we use a hat to indicate empirical measures (e.g., $\hat{P}$).
@@ 2.1 Empirical Optimal Transport
OT [villani2008optimal] is a field of mathematics concerned with the transportation of mass at least effort. We refer readers to [peyre2019computational] for a computational treatment of the subject, and [montesuma2024recent] for applications in machine learning. Let $P,Q\in\mathcal{P}(\Omega)$, and $c:\Omega^{2}\rightarrow\mathbb{R}$ be a ground-cost. OT was originally founded by Monge [monge1781memoire]. In this formulation, we seek for a mapping $T:\Omega\rightarrow\Omega$ such that,
$T_{P\rightarrow Q}^{\star}=\underset{T_{\sharp}P=Q}{\text{arg inf}}\,\int_{\Omega}c(z,T(z))dP(z).$ (2)
@@
Alternatively, Kantorovich [kantorovich1942transfer] proposed a formulation in terms of a transport plan $\gamma\in\Gamma(P,Q)$, where $\Gamma(P,Q)$ is the set of all joint measures with marginals $P$ and $Q$. In this case,
$\gamma^{\star}=\underset{\gamma\in\Gamma(P,Q)}{\text{arg inf}}\,\int_{\Omega}\int_{\Omega}c(z,z^{\prime})\gamma(z,z)$ (3)
Given $p\in[1,+\infty)$, when $c(z,z^{\prime})=d(z,z^{\prime})^{p}$, the previous problems define the $p$-Wasserstein distance,
$\mathbb{W}_{p}(P,Q)^{p}=\underset{\gamma\in\Gamma(P,Q)}{\text{inf}}\,\int_{\Omega}\int_{\Omega}d(z,z^{\prime})^{p}d\gamma(z,z^{\prime}).$ (4)
Equation 4 defines a metric in $\mathcal{P}_{p,ac}(\Omega)$. Conceptually, the Wasserstein distance lifts the metric $d$ on $\Omega$ to $\mathbb{W}_{p}$ on $\mathcal{P}_{p,ac}(\Omega)$. Based on this idea, we can define barycenters and gradient flows in the space of probability measures.
2.2 Gaussian Mixture Optimal Transport
In comparison with equation 1, we can approximate $P$ through a Gaussian Mixture Model (GMM),
$P(z)=\sum_{i=1}^{n}\pi_{i}^{(P)}\mathcal{N}(z|\mu_{i}^{(P)},\Sigma_{i}^{(P)}),$ (5)
where $\theta=\{(\pi_{i}^{(P)},\mu_{i}^{(P)},\Sigma_{i}^{(P)})\}$ are the parameters of the GMM. The general theory of OT between GMMs was first presented in [delon2020wasserstein]. The main advantage is that, when the OT plan is restricted to the set of GMMs, i.e, $\gamma\in\Gamma(P,Q)\cap\text{GMM}_{\infty}(d)$, there is a discrete equivalent formulation in terms of the GMM components,
$\omega^{\star}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}\mathbb{W}_{2}(P_{i},Q_{j})^{2}.$
$\omega^{\star}\in\mathbb{R}^{n\times m}$ is a component-to-component OT plan. The term $\mathbb{W}_{2}(P_{i},Q_{j})$ is the Wasserstein distance between Gaussian components, which has closed-form [takatsu2011wasserstein] in terms of their parameters,
$\mathbb{W}_{2}(P,P^{\prime})^{2}=\lVert\mu-\mu^{\prime}\rVert_{2}^{2}+\text{Tr}(\Sigma+\Sigma^{\prime}-2(\Sigma^{\frac{1}{2}}\Sigma^{\prime}\Sigma^{\frac{1}{2}})^{\frac{1}{2}}).$
Based on these ideas, one may define a Wasserstein-type distance between GMMs,
$\mathbb{MW}_{2}(P,Q)^{2}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{inf}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}\mathbb{W}_{2}(P_{i},Q_{j})^{2}.$ (6)
2.3 Wasserstein Barycenter and its Variants
@@
In the metric setting, the barycenter problem is known as Fréchet [frechet1948elements] or Karcher [karcher1977riemannian] means. In our case, given a finite family of measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$ and a set of barycentric coordinates $\lambda\in\Delta_{K}$, we define the Wasserstein barycenters over $(\mathcal{P}_{p,ac}(\Omega),\mathbb{W}_{p})$ through the following optimization problem,
@@
$P^{\star}=\underset{P\in\mathcal{P}(\Omega)}{\text{arg min}}\,\biggr\{\mathbb{B}_{Q}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(P,Q_{k})^{p}\biggr\}.$ (7)
In general, equation 7 does not have a closed-form solution. However, there are algorithms for computing barycenters for $p=2$, when the measures in $\mathcal{P}$ are either empirical measures [cuturi2014fast], Gaussian measures [agueh2011barycenters, alvarez2016fixed] or Gaussian mixtures [delon2020wasserstein, montesuma2024lighter], and a further link can be made to multi-marginal OT [pass2015multi].
The work of [cuturi2014fast] is of particular interest to us, since theirs was the first algorithm, based on gradient descent, to optimize equation 7 on empirical measures. The idea is to initialize $z_{0,1}^{(P)},\cdots,z_{0,n}^{(P)}$ randomly (e.g., from a Gaussian measure), and iterate
$z_{\tau+1,i}^{(P)}=(1-\alpha)z_{\tau,i}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\gamma_{k}}(z_{\tau,i}^{(P)}),$ (8)
where $T_{\gamma_{k}}(z_{i}^{(P)})=n\sum_{j=1}^{n}\gamma_{k,i,j}z_{j}^{(Q_{k})}$ is the barycentric map between $Q_{k}$ and $P$. This strategy is reminiscent of the fixed-point approach of [alvarez2016fixed]. Furthermore, [montesuma2023multi] shows that, for feature-label joint measures, the labels can be propagated through [redko2019optimal],
$y_{\tau+1,i}^{(P)}=(1-\alpha)y_{\tau,i}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\gamma_{k}}(y_{\tau,i}^{(P)}),$ (9)
where $y_{\tau,i}^{(P)}$ are one-hot encoded vectors.
Previous work has extended equation 8 to the GMM setting, when the components are axis-aligned Gaussians, i.e., $\Sigma_{i}^{(P)}=\text{diag}(\sigma_{i}^{(P)})$. In this case,
$\mu_{i,\tau+1}^{(P)}=(1-\alpha)\mu_{i,\tau}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\omega_{k}}(\mu_{i,\tau}^{(P)}),$ (10)
$\sigma_{i,\tau+1}^{(P)}=(1-\alpha)\sigma_{i,\tau}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\omega_{k}}(\sigma_{i,\tau}^{(P)}),$
There are a few limitations with equations 8 and 10. First, the empirical iterations do not scale well to large datasets. Indeed, they assume the availability of all samples of $Q_{k}$ per iteration. Second, the GMM iterations in equation 10 are restricted to axis-aligned GMM s. This paper addresses these gaps using gradient flows.

### v3

# Wasserstein Gradient Flows for Scalable and Regularized Barycenter Computation

@@ Abstract
Wasserstein barycenters provide a principled approach for aggregating probability measures, while preserving the geometry of their ambient space. Existing discrete methods are not scalable as they assume access to the complete set of samples from the input measures. Meanwhile, neural network approaches do scale well, but rely on complex optimization problems and cannot easily incorporate label information. We address these limitations through gradient flows in the space of probability measures. Through time discretization, we achieve a scalable algorithm that i) relies on mini-batch optimal transport, ii) accepts modular regularization through task-aware functions, and iii) seamlessly integrates supervised information into the ground-cost. We empirically validate our approach on domain adaptation benchmarks that span computer vision, neuroscience, and chemical engineering. Our method establishes a new state-of-the-art barycenter solver, with labeled barycenters consistently outperforming unlabeled ones.
@@ 1 Introduction
Defining the mean or center of a set of probability measures is a fundamental primitive in geometric probability theory [nielsen2020]. Given $K\in\mathbb{N}$ probability measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$ on a metric space $(\Omega,d)$, their Wasserstein barycenter [agueh2011barycenters] is given by,
$P^{\star}=\underset{P\in\mathcal{P}_{2}(\Omega)}{\text{arg min}}\,\biggl\{\mathbb{B}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P,Q_{k})^{2}\biggr\},$ (1)
where $\mathcal{P}_{2}(\Omega)$ is the set of measures over $\Omega$ with finite second moments, and $\lambda=(\lambda_{1},\cdots\lambda_{k})$, $\lambda_{k}\geq 0$, $\forall k$, $\sum_{k}\lambda_{k}=1$ is an array of barycentric coordinates. Equation 1 defines a notion of average of the input measures in $\mathcal{Q}$. The main advantage of these barycenters stems from Optimal Transport (OT) [villani2008optimal], that is, they lift the geometry of the space $(\Omega,d)$ to the space $(\mathcal{P}_{2}(\Omega),\mathbb{W}_{2})$ [kloeckner2010geometric].
Wasserstein barycenters have wide applications to machine learning, such as posterior aggregation in Bayesian inference [srivastava2018scalable], model fusion [singh2020model], ensembling in time series forecasting [coz2023barycenter], data augmentation [zhu2023interpolation], and Domain Adaptation (DA) [montesuma2021wasserstein, montesuma2023multi, montesuma2024lighter]. In each application, the main advantage is geometric fidelity, that is, the Wasserstein barycenter captures and preserves the underlying structure of samples. Figure 1 shows a conceptual illustration.
(a) Conceptual Illustration.
(b) $\mathbb{B}$.
(c) $\mathbb{B}+\mathbb{V}$.
(d) $\mathbb{B}+\mathbb{V}+\mathbb{U}$.
Figure 1: In (a), we show an illustration of our flow. We show the initial barycenter measure $P_{0}$ which flow towards the solution $P^{\star}$. In red, green, and orange, we show the input measures’ mini-batches as solid dots. The flow is given by the combination of gradients $v_{\tau,i}=\sum_{k=1}^{K}\lambda_{k}\nabla\mathbb{W}_{2}(P_{\tau},Q_{k})$. In parallel, (b) shows a pathological barycenter calculation, with fuzzy labels and unclear class boundaries. Adding regularizers $\mathbb{V}$ (b) and $\mathbb{U}$ (c) progressively resolves this issue.
Despite its advantages, Wasserstein barycenters remain challenging beyond the Gaussian setting [altschuler2022wasserstein]. We have identified three gaps in the current literature. First, the seminal algorithm of cuturi2014fast requires having access to the complete set of input measures’ samples at once, making it intractable for large datasets. Neural network methods [fan2020scalable, korotin2021continuous, kolesov2024estimating, gazdieva2024robust] alleviate this issue by operating on mini-batches, but parametrize the barycentric measure with $\mathcal{O}(K)$ neural networks, coupling model complexity with the number of input measures. Second, neural methods cannot seamlessly incorporate label information into their ground-cost, limiting their performance in supervised tasks (cf. Table 2). Third, the objective function in equation 1 only accounts for distributional fit. In practice, the barycentric measure must satisfy additional structural properties, such as class separation. Current barycenter solvers do not offer a principled way of enforcing these properties.
Recent methods study the barycenter problem through the gradient flow perspective. First, chewi2020gradient established a gradient flow algorithm on the Bures-Wasserstein manifold, i.e., on Gaussian measures. Second, chizat2025doubly describes the doubly regularized barycenter problem through the same lens. Here, in addition to entropic regularization at the level of OT, they proposed an outer regularization that penalizes the entropy of the barycentric measure, which leads to a Noisy Particle Gradient Descent (NPGD) algorithm for empirical measures. Although this algorithm incorporates outer regularization, it has the same limitation as cuturi2014fast, as it assumes full-batch access to the samples of input measures.
We address the aforementioned challenges through the lens of gradient flows in the space of probability measures [ambrosio2008gradient, santambrogio2017euclidean]. We therefore conceptualize the barycenter problem as the flow from an initial measure $P_{0}=\mathcal{N}(0,\text{Id})$ following the Wasserstein gradient of the functional,
$\mathbb{F}(P)=\mathbb{B}(P)+\mathbb{R}(P),$ (2)
where $\mathbb{B}$ denotes the barycenter functional (cf. equation 1), and $\mathbb{R}$ denotes additional regularizing functionals (see Sections 2.3 and 3.3, and Appendix C.2).
Following the established gradient flow literature [santambrogio2017euclidean], we decompose $\mathbb{R}$ into internal, potential, and interaction energies, denoted $\mathbb{G}$, $\mathbb{V}$, and $\mathbb{U}$, respectively. This decomposition generalizes the double regularization strategy of chizat2025doubly. Indeed, setting $\mathbb{V}=\mathbb{U}=0$ and $\mathbb{G}(P)=\eta_{G}H(P)$, where $H(P)$ denotes the entropy of $P$, corresponds to their exact framework. Similarly, alvarez2021dataset devised an approach for flowing a single dataset towards a known target. Our setting differs structurally from that. First, we go beyond the single measure objective, optimizing the multi-measure barycentric objective (cf. equation 1). Second, while alvarez2021dataset transforms an existing dataset towards a known target, ours synthesizes a measure (i.e., the barycentric measure) from noise.
Our contributions are as follows. (i) A mini-batch, time discretized gradient flow algorithm for regularized empirical Wasserstein barycenters (Algorithm 1). By randomly sampling the input measures, we reduce the computational complexity of traditional barycenter methods [cuturi2014fast, montesuma2023multi, chizat2025doubly] achieving $2\times$ to $50\times$ speedups with respect to discrete solvers (Figure 5). (ii) Modular, task-aware regularizing functionals for Wasserstein barycenters (Section 3.3). Following previous gradient flow literature [santambrogio2017euclidean, alvarez2021dataset], we decompose $\mathbb{R}$ into internal, potential, and interaction energies, allowing incorporation of plug-and-play regularizers into the barycenter problem. This goes beyond chizat2025doubly, who considered a particular internal energy as an outer regularization strategy. (iii) Extensive DA benchmarking (Section 4). We empirically show that our algorithm produces barycenters that respect class structure, by incorporating both appropriate functionals (Section 3.3), and labels into the ground-cost of optimal transport (Section 3.2). On 5 DA benchmarks, spanning computer vision, neuroscience, and chemical engineering, we establish a comparison between empirical and neural barycenter solvers, showing that structured costs are essential for DA performance.
Paper organization. Section 2 reviews OT, Wasserstein barycenters and gradient flows. Section 3 develops our empirical, time-discretized gradient flow framework and algorithms. Section 4 presents our experiments. Finally, Section 5 concludes this paper.
@@ 2 Background
Throughout this paper $(\Omega,d)$ is a metric space, and its elements are denoted $z\in\Omega$. We denote $\mathcal{P}_{2}(\Omega)$ as the set of measures with finite second moments [villani2008optimal, Definition 6.4]. We work with empirical measures,
$\text{Emp}_{n}(\Omega)=\biggr\{\hat{P}\in\mathcal{P}_{2}(\Omega):\hat{P}=\frac{1}{n}\sum_{i=1}^{n}\delta_{z_{i}^{(P)}}\biggr\},$ (3)
where $\delta_{z_{0}}(z)=\delta(z-z_{0})$ denotes the Dirac measure centered at $z_{0}\in\Omega$. Given a measure $P\in\mathcal{P}_{2}(\Omega)$, we denote by $\hat{P}\in\text{Emp}_{n}(\Omega)$ its empirical approximation based on i.i.d. samples $z_{1}^{(P)},\cdots,z_{n}^{(P)}\sim P$. The superscript $\cdot^{(P)}$ indicates the measure from which the samples originate.
2.1 Optimal Transport
@@
OT is a field of mathematics [villani2008optimal] concerned with mass transportation at least effort [peyre2019computational, montesuma2024recent]. It was originally founded by monge1781memoire, and seeks a mapping such that,
$T_{P\rightarrow Q}^{\star}=\underset{T_{\sharp}P=Q}{\text{arg inf}}\,\mathbb{E}_{z\sim P}[d(z,T(z))^{2}].$ (4)
@@
$T_{P\rightarrow Q}^{\star}$ is an OT mapping between $P$ and $Q$, and $T_{\sharp}$ is the pushforward mapping $P$ to $Q$. Alternatively, kantorovich1942transfer proposed a formulation in terms of a transport plan $\gamma\in\Gamma(P,Q)$, where $\Gamma(P,Q)$ is the set of all joint measures with marginals $P$ and $Q$. In this case,
$\gamma^{\star}=\underset{\gamma\in\Gamma(P,Q)}{\text{arg inf}}\,\mathbb{E}_{(z,z^{\prime})\sim\gamma}[d(z,z^{\prime})^{2}],$ (5)
where $\gamma^{\star}$ denotes an OT plan. OT defines the so-called 22-Wasserstein distance, through the infimum value of the optimization problem in equations 4 and 5,
$\mathbb{W}_{2}(P,Q)^{2}=\underset{\gamma\in\Gamma(P,Q)}{\text{inf}}\,\mathbb{E}_{(z,z^{\prime})\sim\gamma}[d(z,z^{\prime})^{2}].$ (6)
Equation 6 defines a metric in $\mathcal{P}_{2}(\Omega)$ which bridges the notions of Euclidean (i.e., center of mass) and Wasserstein barycenters. An alternative consists of regularizing the OT problem with the entropy of $\gamma$, resulting in the Sinkhorn divergence cuturi2013sinkhorn,
$\mathbb{W}_{2,\epsilon}(P,Q)^{2}=\underset{\gamma\in\Gamma(P,Q)}{\text{inf}}\,\mathbb{E}[d(z,z^{\prime})^{2}]+\lambda H(\gamma),$ (7)
where $H(\gamma)=\text{KL}(\gamma|P\otimes Q)$, for $(P\otimes Q)(z)=P(z)Q(z)$, and KL is the Kullback Leibler divergence. Equation 7 enjoys many desirable properties. First, by relying on the matrix scaling algorithm of sinkhorn1964relationship, it has complexity $\mathcal{O}(n^{2})$ per iteration. Second, it can be vectorized in modern GPUs, leading to dramatic speed ups. Third, since $H(\gamma)$ is strictly convex, the overall optimization problem in equation 7 is convex, leading to a unique solution.
2.2 Wasserstein Barycenter
@@
In the metric setting, the barycenter problem is known as frechet1948elements or karcher1977riemannian means. In our case, given a finite family of measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$ and a set of barycentric coordinates $\lambda\in\Delta_{K}=\{a\in\mathbb{R}^{K}_{\geq 0}:\sum_{k}a_{k}=1\}$, we define the Wasserstein barycenter over $(\mathcal{P}_{2}(\Omega),\mathbb{W}_{2})$ through the following optimization problem,
@@
$\hat{P}^{\star}=\underset{\hat{P}\in\text{Emp}_{n}(\Omega)}{\text{arg min}}\,\biggr\{\mathbb{B}_{2,\epsilon}(\hat{P}|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2,\epsilon}(\hat{P},\hat{Q}_{k})^{2}\biggr\}.$ (8)
[…]